1. Simon recently received a credit card with an 17% nominal interest rate. With the card, he purchased an Apple iPhone 5 for $440. The minimum payment on the card is only $20 per month.

  1. If Simon makes the minimum monthly payment and makes no other charges, how many months will it be before he pays off the card? Do not round intermediate calculations. Round your answer to the nearest month.
    ____ month(s)

  2. If Simon makes monthly payments of $60, how many months will it be before he pays off the debt? Do not round intermediate calculations. Round your answer to the nearest month.
    ___ month(s)

  3. How much more in total payments will Simon make under the $20-a-month plan than under the $60-a-month plan. Do not round intermediate calculations. Round your answer to the nearest cent.
    $____

2. Find the amount to which $500 will grow under each of these conditions:

  1. 14% compounded annually for 4 years. Do not round intermediate calculations. Round your answer to the nearest cent.
    $ ____

  2. 14% compounded semiannually for 4 years. Do not round intermediate calculations. Round your answer to the nearest cent.
    $ _____

  3. 14% compounded quarterly for 4 years. Do not round intermediate calculations. Round your answer to the nearest cent.
    $ _____

  4. 14% compounded monthly for 4 years. Do not round intermediate calculations. Round your answer to the nearest cent.
    $ ______

  5. 14% compounded daily for 4 years. Do not round intermediate calculations. Round your answer to the nearest cent.
    $ _______
Answer #1

You have asked multiple unrelated questions in the same post. Not only that, each question has multiple sub parts. I have addressed the first one. Please post the balance questions separately one by one.

1. Simon recently received a credit card with an 17% nominal interest rate. With the card, he purchased an Apple iPhone 5 for $440. The minimum payment on the card is only $20 per month.

If Simon makes the minimum monthly payment and makes no other charges, how many months will it be before he pays off the card? Do not round intermediate calculations. Round your answer to the nearest month.

=NPER(Rate, PMT, PV, FV) = NPER(17%/12,20,-440,0) = 26.55 = 27 month(s)

If Simon makes monthly payments of $60, how many months will it be before he pays off the debt? Do not round intermediate calculations. Round your answer to the nearest month.

=NPER(Rate, PMT, PV, FV) = NPER(17%/12,60,-440,0) = 7.80 =  8 month(s)

How much more in total payments will Simon make under the $20-a-month plan than under the $60-a-month plan. Do not round intermediate calculations. Round your answer to the nearest cent.

20 x 26.55 - 60 x 7.80 = $ 63.14